OpenStudy (triarimig3):

The height of a cone is two times its base diameter. What is the volume of the cone in terms of its base radius r?

rebeccaxhawaii (rebeccaxhawaii):

WELCOME TO OS !!!! so are you just writing an expression?

OpenStudy (welshfella):

the volume of a cone is given by (1/3) *pi*r^2* h where r = radius of the base and h = the height

OpenStudy (welshfella):

diameter = twice the radius so if the height is twice the diameter what is the relation between the height and the radius? can you figure that out?

OpenStudy (triarimig3):

No, I'm doing this plato learning thing for credit recovery, I do not understand it at all.

OpenStudy (welshfella):

OK lets call the diameter d so d= 2r and h = 2d therefore h = 2(2r) = 4r

OpenStudy (welshfella):

the height is 4 times the radius of the base OK?

OpenStudy (triarimig3):

Ok, I understand it better, it's just that Geometry is not my thing.

OpenStudy (welshfella):

so if we replace the h in the formula for the volume by 4r we'll have what's required Vol = (1/3) pi r^2 * h = (1/3) pi r^2 * 4r = (1/3) pi r^3 * 4 = (4/3) pi r^3

OpenStudy (triarimig3):

Thank you.

OpenStudy (welshfella):

yw